Can you help me solve this problem and show the steps?The local transformer shown below provides the last voltage stepdown in a power distribution system. A common sight on utilit
Answer #1

Turns ratio = N2/N1= V2/V1= 240/13800= 1/57.5

n=57.5

1:n=1:57.5

Total secondary current= I2 = 10*200 = 2000A

(Since there are 10 houses which takes 200A each)

Power rating of transformer= V2*I2

= 240*2000

=480 KVA

Primary current I1= 480kva/13.8kv

Therefore , I1 = 34.782A